Sally and Garry's combined age is 39,27 years more than the combined age of Garry and Nora,which is 25 years less than Sally and Nora's combined age.In how many years'time will the sum of all their ages be 188.
Answer by ankor@dixie-net.com(22740)

Sally and Garry's combined age is 39, 27 years more than the combined age of Garry and Nora, which is 25 years less than Sally and Nora's combined age.
In how many years time will the sum of all their ages be 188?
:
"Sally and Garry's combined age is 39"
S + G = 39
:
"27 years more than the combined age of Garry and Nora,"
N + G = 39-27
N + G = 12
:
"which is 25 years less than Sally and Nora's combined age."
N + G = S + N - 25
subtract N from both sides
G = S - 25
:
Using the first equation, replace G with (S-25)
S + (S-25) = 39
2S = 39 + 25
2S = 64
S = 64/2
S = 32 is Sally's age
then
G + 32 = 39
G = 39 - 32
G = 7 yrs is Gary's age
:
Find N
N + G = 12
N + 7 = 12
N = 12 - 7
N = 5 yrs is Nora's age
:
:
Check solutions in the last equation
N + G = S + N - 25
5 + 7 = 32 + 5 - 25
12 = 37 - 25; confirms our solution of: G=7, S=32, N=5
:
"In how many years time will the sum of all their ages be 188?"
let y = no. of yrs for this to be true
(y+7) + (y+32) + (y+5) = 188
3y + 44 = 188
3y = 188 - 44
3y = 144
y = 144/3
y = 48 yrs
:
:
Check this:
(y+7) + (y+32) + (y+5) =
48+7 + 48+32 + 48+5 =
55 + 80 + 53 = 188